OpenStudy (anonymous):

determine whether the lines are parallel perpendicular or neither -10x-5y=2 8y=-4x+12

OpenStudy (anonymous):

first try to get them into y=mx+b form -10x-5y=2 +10x +10x -5y=10x+2 divide by -5 y=-2x-(2/5) 8y=-4x+12 divide by 8 y=-(1/2)x+(3/2) now look at the m-values -2 and -1/2 what do you think the answer is?

OpenStudy (anonymous):

they are opposite and reciprocal so it is perpendicular right?

OpenStudy (anonymous):

no they are not opposite they are both negative so it would be neither

OpenStudy (anonymous):

oh ok thank you so much!!!! i appreciate it!!!

OpenStudy (anonymous):

yep